S. Would 0/. Coh Eidnce interal btnaue of wic eps invlved in Cheir teshn
Answer #1

(A).

99.94% confidence intervalis wider than 90% confidence interval .

Consider z test

For 90% confidence, \alpha=1-.90=0.10,\alpha​​​​​​/2 =0.05

Z\alpha/2 =1.96

For 99.94% confidence,\alpha=1-0.994=0.006, \alpha/2=0.003

Z\alpha​​​​​​/2=2.748

Confidence interval

X+-Z\alpha/2(\sigma/√n)

Hence,after the calculation we can show that 99.94% confidence interval is more wider than 90% confidence interval.

(B) . Null hypothesis Ho: \mu =\muo

Alternative hypothesis Ha:\mu< \muo or \mu>\muo (for one tailed test)

Null hypothesis Ho :\mu ,=\muo

Alternative hypothesis Ha :\mu\neq\muo (for two tailed test).

Test statistic Z=(x- \muo)/(\sigma/√n)

P value =1-p(z). For one tailed test

P value =2(1-p(z)) for two tailed test.

Reject Ho if p Value <\alpha .

Not reject Ho if p value>\alpha.